OpenStudy (cutiecomittee123):

which of the following could be an example of a function with the domain [a, infinity) and range [b, infinity) where a>0 and b>0 f(x)= square root of x+a -b outside the root sign. f(x)= square root of x-a +b outside the square root sign. f(x)= the cube root of (x-b) +a outside of the root sign f(x)=the cube root of x+a -b outside of the root sign.

geerky42 (geerky42):

\(f(x) = \sqrt x\) has domain \([0,\infty)\) and range \([0,\infty)\)

geerky42 (geerky42):

So that would left us either A or B.

OpenStudy (cutiecomittee123):

Okay